# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 2.2 Hopf algebroid structure
Define two moment maps $\mu_{l},\mu_{r}:{\mathbb{F}}[[p]]\to{\cal U}$ by
@@
Let $\gamma_{(1)}=\gamma\widetilde{\otimes}1,\gamma_{(2)}=1\widetilde{\otimes}\gamma$ and $p^{*}_{(i)}=p\gamma^{-2}_{(i)}$ $(i=1,2)$. Let us define two algebra homomorphisms $\Delta^{D}:{\cal U}\to{\cal U}\;\widetilde{\otimes}\;{\cal U}$ and $\varepsilon^{D}:{\cal U}\to{\mathbb{C}}$ by
$\Delta^{D}(\gamma^{\pm 1/2})=\gamma^{\pm 1/2}\widetilde{\otimes}\gamma^{\pm 1/2},$ (2.16)
$\Delta^{D}(\psi^{\pm}(z))=\psi^{\pm}(\gamma^{\mp 1/2}_{(2)}z)\widetilde{\otimes}\psi^{\pm}(\gamma^{\pm 1/2}_{(1)}z),$ (2.17)
$\Delta^{D}(x^{+}(z))=1\widetilde{\otimes}x^{+}(\gamma^{-1/2}_{(1)}z)+x^{+}(\gamma^{1/2}_{(2)}z)\widetilde{\otimes}\psi^{-}(\gamma^{-1/2}_{(1)}z),$ (2.18)
$\Delta^{D}(x^{-}(z))=x^{-}(\gamma^{-1/2}_{(2)}z)\widetilde{\otimes}1+\psi^{+}(\gamma^{-1/2}_{(2)}z)\widetilde{\otimes}x^{-}(\gamma^{1/2}_{(1)}z),$ (2.19)
$\Delta^{D}(\mu_{l}(F(\mathfrak{z},p)))=\mu_{l}(F(\mathfrak{z},p))\widetilde{\otimes}1,\qquad\Delta^{D}(\mu_{r}(F(\mathfrak{z},p)))=1\widetilde{\otimes}\mu_{r}(F(\mathfrak{z},p)),$ (2.20)
$\varepsilon^{D}(\gamma^{1/2})=\varepsilon^{D}(C)=1,\qquad\varepsilon^{D}(\psi^{\pm}(z))=1,\qquad\varepsilon^{D}(x^{\pm}(z))=0,$ (2.21)
$\varepsilon^{D}(\mu_{l}(F(\mathfrak{z},p)))=\varepsilon^{D}(\mu_{r}(F(\mathfrak{z},p)))=F(\mathfrak{z},p).$ (2.22)
@@
$(\Delta^{D}\widetilde{\otimes}{\rm id})\circ\Delta^{D}=({\rm id}\widetilde{\otimes}\Delta^{D})\circ\Delta^{D},$ (2.23)
$(\varepsilon^{D}\widetilde{\otimes}{\rm id})\circ\Delta^{D}={\rm id}=({\rm id}\widetilde{\otimes}\varepsilon^{D})\circ\Delta^{D}.$ (2.24)
We also define an algebra anti-homomorphism $S^{D}:{\cal U}\to{\cal U}$ by
@@
$m\circ({\rm id}\widetilde{\otimes}S^{D})\circ\Delta^{D}(a)=\mu_{l}(\varepsilon^{D}(a)1)\qquad\forall a\in{\cal U},$
@@
These Propositions indicate that $({\cal U},\Delta^{D},\varepsilon^{D},\mu_{l},\mu_{r},S^{D})$ is a Hopf algebroid[6, 21, 25, 26].
@@
2.3 Representations of ${U}_{t_{1},t_{2},p}({{\mathfrak{gl}}}_{1,tor})$
Let ${\cal V}$ be a ${\cal U}$-module. For $(k,l)\in{\mathbb{C}}^{2}$, we say that ${\cal V}$ has level $(k,l)$, if the central elements $\gamma$ and $C$ act as
$\gamma\cdot\xi=\hbar^{k/2}\xi,\qquad C\cdot\xi=\hbar^{l/2}\xi\qquad\forall\xi\in{\cal V}.$
@@ 2.3 Representations of ${U}_{t_{1},t_{2},p}({{\mathfrak{gl}}}_{1,tor})$
2.3.1 The level-$(1,N)$ representation
Let us introduce one more Heisenberg algebra generated by $h,\alpha$ with
$[h,\alpha]=1.$
@@ 2.3.1 The level-$(1,N)$ representation
${|0\rangle}^{(1,N)}_{v}:=v^{\alpha}e^{\Lambda_{0}}e^{N\Lambda_{0}^{\perp}}1.$ (2.25)
@@
$e^{\pm h}x^{\pm c^{\perp}}\cdot{|0\rangle}^{(1,N)}_{v}=v^{\pm 1}x^{\pm N}{|0\rangle}_{v}^{(1,N)},\quad\gamma\cdot{|0\rangle}^{(1,N)}_{v}=\hbar^{1/2}{|0\rangle}^{(1,N)}_{v},\quad C^{\pm 1}\cdot{|0\rangle}^{(1,N)}_{v}=\hbar^{\pm N/2}{|0\rangle}^{(1,N)}_{v}.$
Let ~$\widetilde{{\cal F}}^{(1,N)}_{v}={\mathbb{C}}[\alpha_{-m}\ (m>0)]{|0\rangle}^{(1,N)}_{v}$ be a Fock space on which the Heisenberg subalgebra $\{\alpha_{m}\ (m\in{\mathbb{Z}}_{\not=0})\}$ acts as
@@
for $m>0$, $\xi\in\widetilde{{\cal F}}^{(1,N)}_{v}$. Note that $p^{*}=p\hbar^{-1}$ on ~$\widetilde{{\cal F}}^{(1,N)}_{v}$. We set
${\cal F}^{(1,N)}_{v}=\bigoplus_{m\in{\mathbb{Z}}}\widetilde{{\cal F}}^{(1,N)}_{v}\otimes e^{mQ}.$
@@
The following assignment gives a level $(1,N)$ representation of ${\cal U}$ on ${\cal F}^{(1,N)}_{v}$.
$x^{+}(z)=e^{h}(z^{-1}\hbar^{1/2})^{{c^{\perp}}}\exp\left\{-\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha_{-n}z^{n}\right\}\exp\left\{\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha_{n}z^{-n}\right\},$ (2.26)
$x^{-}(z)=e^{-h}(z^{-1}\hbar^{1/2})^{-{c^{\perp}}}\exp\left\{\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha^{\prime}_{-n}z^{n}\right\}\exp\left\{-\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha^{\prime}_{n}z^{-n}\right\},$ (2.27)
$\psi^{+}(\hbar^{1/4}z)=\hbar^{-{c^{\perp}}/2}\exp\left\{-\sum_{n>0}\frac{p^{n}}{1-p^{n}}\alpha_{-n}z^{n}\right\}\exp\left\{\sum_{n>0}\frac{1}{1-p^{n}}\alpha_{n}z^{-n}\right\},$ (2.28)
$\psi^{-}(\hbar^{-1/4}z)=\hbar^{{c^{\perp}}/2}\exp\left\{-\sum_{n>0}\frac{1}{1-p^{n}}\alpha_{-n}z^{n}\right\}\exp\left\{\sum_{n>0}\frac{p^{n}}{1-p^{n}}\alpha_{n}z^{-n}\right\}.$ (2.29)
@@
2.3.2 The level-(0,1) representation
We next consider the level-(0,1) representation of ${\cal U}$ called the $q$-Fock representation.
For $u\in{\mathbb{C}}^{\times}$, let ${\cal F}_{u}$ be a vector space spanned by ${|\lambda\rangle}_{u}$ $(\lambda\in{\cal P})$, where
${\cal P}=\{\lambda=(\lambda_{1},\lambda_{2},\cdots)\ |\ \lambda_{i}\geq\lambda_{i+1},\ \lambda_{i}\in{\mathbb{Z}}_{\geq 0},\ \lambda_{l}=0\ \mbox{for sufficiently large}\ l\ \}.$ (2.30)
We denote by $\ell(\lambda)$ the length of $\lambda\in{\cal P}$ i.e. $\lambda_{\ell(\lambda)}>0$ and $\lambda_{\ell(\lambda)+1}=0$. We also set $|\lambda|=\sum_{i\geq 1}\lambda_{i}$ and denote by $\lambda^{\prime}$ the conjugate of $\lambda$.
@@ 2.3.2 The level-(0,1) representation
The following action gives a level (0,1) representation of ${\cal U}$ on ${\cal F}_{u}$. We denote this by ${\cal F}^{(0,1)}_{u}$.
$\gamma^{1/2}{|\lambda\rangle}_{u}={|\lambda\rangle}_{u},$ (2.31)
$x^{+}(z){|\lambda\rangle}_{u}=a^{+}(p)\sum_{i=1}^{\ell(\lambda)+1}A^{+}_{\lambda,i}(p)\delta(u_{i}/z){|\lambda+{\bf 1}_{i}\rangle}_{u},$ (2.32)
$x^{-}(z){|\lambda\rangle}_{u}=a^{-}(p)\sum_{i=1}^{\ell(\lambda)}A^{-}_{\lambda,i}(p)\delta(t_{1}u_{i}/z){|\lambda-{\bf 1}_{i}\rangle}_{u},$ (2.33)
$\psi^{+}(z){|\lambda\rangle}_{u}=B^{+}_{\lambda}(u/z;p){|\lambda\rangle}_{u},$ (2.34)
$\psi^{-}(z){|\lambda\rangle}_{u}=B^{-}_{\lambda}(z/u;p){|\lambda\rangle}_{u},\quad$ (2.35)
@@
$a^{+}(p)$ $=$ $(1-t)\frac{(p\hbar;p)_{\infty}(p/t_{2};p)_{\infty}}{(p;p)_{\infty}(p/q;p)_{\infty}},$ (2.36)
$a^{-}(p)$ $=$ $(1-t^{-1})\frac{(p/\hbar;p)_{\infty}(pt_{2};p)_{\infty}}{(p;p)_{\infty}(pq;p)_{\infty}},$ (2.37)
$A^{+}_{\lambda,i}(p)$ $=$ $\prod_{j=1}^{i-1}\frac{\theta(t_{2}u_{i}/u_{j})\theta(\hbar^{-1}u_{i}/u_{j})}{\theta(t_{1}^{-1}u_{i}/u_{j})\theta(u_{i}/u_{j})},$ (2.38)
$A^{-}_{\lambda,i}(p)$ $=$ $\prod_{j=i+1}^{\ell(\lambda)}\frac{\theta(\hbar^{-1}u_{j}/u_{i})}{\theta(u_{j}/u_{i})}\prod_{j=i+1}^{\ell(\lambda)+1}\frac{\theta(t_{2}u_{j}/u_{i})}{\theta(t_{1}^{-1}u_{j}/u_{i})},$ (2.39)
$B^{+}_{\lambda}(u/z;p)$ $=$ $\prod_{i=1}^{\ell(\lambda)}\frac{\theta(t_{2}^{-1}u_{i}/z)}{\theta(t_{1}u_{i}/z)}\prod_{i=1}^{\ell(\lambda)+1}\frac{\theta(\hbar u_{i}/z)}{\theta(u_{i}/z)},$ (2.40)
$B^{-}_{\lambda}(z/u;p)$ $=$ $\prod_{i=1}^{\ell(\lambda)}\frac{\theta(t_{2}z/u_{i})}{\theta(t_{1}^{-1}z/u_{i})}\prod_{i=1}^{\ell(\lambda)+1}\frac{\theta(\hbar^{-1}z/u_{i})}{\theta(z/u_{i})}.$ (2.41)
@@
Remark. In [32] it is conjectured that the level-(0,1) representation in Theorem 2.6 gives a geometric action of ${\cal U}$ on the equivariant elliptic cohomology of the Hilbert schemes $\bigoplus_{n}\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$ by identifying ${|\lambda\rangle}_{u}$ with the fixed point class $[\lambda]$ in $\bigoplus_{n}\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$. The class $[\lambda]$ with $|\lambda|=n$ should be realized in terms of the elliptic stable envelopes $\mathrm{Stab}_{\mathfrak{C}}$ for $\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$ constructed in [44] (see Sec.3.2 below) by
$[\lambda]=\sum_{\mu\in{\cal P}\atop|\mu|=n}\mathrm{Stab}^{-1}_{\mathfrak{C}}(\mu)|_{\lambda}\mathrm{Stab}_{\mathfrak{C}}(\mu)$
@@
3 Elliptic Stable Envelopes for $\mathrm{E}_{T}({\cal M}(n,r))$
In this section we summarize basic facts on the instanton moduli space ${\cal M}(n,r)$ and the elliptic stable envelopes for the equivariant elliptic cohomology $\mathrm{E}_{T}({\cal M}(n,r))$ following [1, 44, 4].
3.1 The instanton moduli space ${\cal M}(n,r)$
Let ${\cal M}(n,r)$ be the moduli space of framed rank $r$ torsion free sheaves ${\cal S}$ on ${\mathbb{P}}^{2}$ with $c_{2}({\cal S})=n$. A framing of a sheaf ${\cal S}$ means a choice of isomorphism
${\cal S}|_{{\mathbb{P}}}\ \to\ {\cal O}^{\oplus r}_{L_{\infty}},$
where $L_{\infty}\subset{\mathbb{P}}^{2}$ is chosen as a line at infinity of a plane in ${\mathbb{P}}^{2}$. The space ${\cal M}(n,r)$ is isomorphic to the Nakajima quiver variety associated to the quiver consisting of one vertex and one loop with the vertex dimension $n$ and the framing dimension $r$, and known as the Atiyah-Drinfeld-Hitchin-Manin (ADHM) instanton moduli space. One has a natural action of $G=GL(r)\times GL(2)$ on ${\cal M}(n,r)$. Let $T$ be the maximal torus of $G$ and set

### v2

@@ 2.2. Hopf algebroid structure
Define two moment maps $\mu_{l},\mu_{r}:{\mathbb{F}}[[p]]\to{\mathcal{U}}$ by
@@
Let $\gamma_{(1)}=\gamma\widetilde{\otimes}1,\gamma_{(2)}=1\widetilde{\otimes}\gamma$ and $p^{*}_{(i)}=p\gamma^{-2}_{(i)}$ $(i=1,2)$. Let us define two algebra homomorphisms $\Delta^{D}:{\mathcal{U}}\to{\mathcal{U}}\;\widetilde{\otimes}\;{\mathcal{U}}$ and $\varepsilon^{D}:{\mathcal{U}}\to{\mathbb{C}}$ by
$\Delta^{D}(\gamma^{\pm 1/2})=\gamma^{\pm 1/2}\widetilde{\otimes}\gamma^{\pm 1/2},$
$\Delta^{D}(\psi^{\pm}(z))=\psi^{\pm}(\gamma^{\mp 1/2}_{(2)}z)\widetilde{\otimes}\psi^{\pm}(\gamma^{\pm 1/2}_{(1)}z),$
$\Delta^{D}(x^{+}(z))=1\widetilde{\otimes}x^{+}(\gamma^{-1/2}_{(1)}z)+x^{+}(\gamma^{1/2}_{(2)}z)\widetilde{\otimes}\psi^{-}(\gamma^{-1/2}_{(1)}z),$
$\Delta^{D}(x^{-}(z))=x^{-}(\gamma^{-1/2}_{(2)}z)\widetilde{\otimes}1+\psi^{+}(\gamma^{-1/2}_{(2)}z)\widetilde{\otimes}x^{-}(\gamma^{1/2}_{(1)}z),$
$\hskip-19.91684pt\Delta^{D}(\mu_{l}(F(\mathfrak{z},p)))=\mu_{l}(F(\mathfrak{z},p))\widetilde{\otimes}1,\ \Delta^{D}(\mu_{r}(F(\mathfrak{z},p)))=1\widetilde{\otimes}\mu_{r}(F(\mathfrak{z},p)),$
$\varepsilon^{D}(\gamma^{1/2})=\varepsilon^{D}(C)=1,\ \varepsilon^{D}(\psi^{\pm}(z))=1,\ \varepsilon^{D}(x^{\pm}(z))=0,$
$\varepsilon^{D}(\mu_{l}(F(\mathfrak{z},p)))=\varepsilon^{D}(\mu_{r}(F(\mathfrak{z},p)))=F(\mathfrak{z},p).$
@@
$(\Delta^{D}\widetilde{\otimes}{\rm id})\circ\Delta^{D}=({\rm id}\widetilde{\otimes}\Delta^{D})\circ\Delta^{D},$
$(\varepsilon^{D}\widetilde{\otimes}{\rm id})\circ\Delta^{D}={\rm id}=({\rm id}\widetilde{\otimes}\varepsilon^{D})\circ\Delta^{D}.$
We also define an algebra anti-homomorphism $S^{D}:{\mathcal{U}}\to{\mathcal{U}}$ by
@@
$m\circ({\rm id}\widetilde{\otimes}S^{D})\circ\Delta^{D}(a)=\mu_{l}(\varepsilon^{D}(a)1)\qquad\forall a\in{\mathcal{U}},$
@@
These Propositions indicate that $({\mathcal{U}},\Delta^{D},\varepsilon^{D},\mu_{l},\mu_{r},S^{D})$ is a Hopf algebroid[6, 21, 25, 26].
@@
2.3. The level-$(1,N)$ representation
Definition 2.2.
Let $\mathcal{V}$} be a $\mathcal{U}$}-module. For $(k,l)\in{\mathbb{C}}^{2}$, we say that $\mathcal{V}$} has level $(k,l)$, if the central elements $\gamma$ and $C$ act as [^2] We changed the definition of the level of representation from the one given in [32] so that our level $(k,l)$ is the level (k,l)(k,-l) there.
$\gamma\cdot\xi=\hbar^{k/2}\xi,\qquad C\cdot\xi=\hbar^{l/2}\xi\qquad\forall\xi\in{\mathcal{V}}.$
@@ 2.3. The level-$(1,N)$ representation
${|0\rangle}^{(1,N)}_{v}:=v^{\alpha}e^{\Lambda^{\perp}_{0}}e^{N\Lambda_{0}}1.$ (2.7)
@@
$e^{\pm h}x^{\pm c}\cdot{|0\rangle}^{(1,N)}_{v}=v^{\pm 1}x^{\pm N}{|0\rangle}_{v}^{(1,N)},\quad\gamma\cdot{|0\rangle}^{(1,N)}_{v}=\hbar^{1/2}{|0\rangle}^{(1,N)}_{v}$$\quad$
$C\cdot{|0\rangle}^{(1,N)}_{v}=\hbar^{N/2}{|0\rangle}^{(1,N)}_{v}$.
Let $\widetilde{{\mathcal{F}}}^{(1,N)}_{v}={\mathbb{C}}[\alpha_{-m}\ (m>0)]{|0\rangle}^{(1,N)}_{v}$ be a Fock space on which the Heisenberg subalgebra $\{\alpha_{m}\ (m\in{\mathbb{Z}}_{\not=0})\}$ acts as
@@
for $m>0$, $\xi\in\widetilde{{\mathcal{F}}}^{(1,N)}_{v}$. Note that $p^{*}=p\hbar^{-1}$ on $\widetilde{{\mathcal{F}}}^{(1,N)}_{v}$. We set
${\mathcal{F}}^{(1,N)}_{v}=\bigoplus_{m\in{\mathbb{Z}}}\widetilde{{\mathcal{F}}}^{(1,N)}_{v}\otimes e^{mQ}.$
@@
The following assignment gives a level-$(1,N)$ representation of $\mathcal{U}$} on ${\mathcal{F}}^{(1,N)}_{v}$.
$\hskip-28.45274ptx^{+}(z)=e^{h}(z^{-1}\hbar^{1/2})^{-c}\exp\left\{-\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha_{-n}z^{n}\right\}\exp\left\{\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha_{n}z^{-n}\right\},$
$\hskip-28.45274ptx^{-}(z)=e^{-h}(z^{-1}\hbar^{1/2})^{c}\exp\left\{\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha^{\prime}_{-n}z^{n}\right\}\exp\left\{-\sum_{n>0}\frac{\hbar^{n/2}}{1-\hbar^{n}}\alpha^{\prime}_{n}z^{-n}\right\},$
$\hskip-28.45274pt\psi^{+}(\hbar^{1/4}z)=\hbar^{c/2}\exp\left\{-\sum_{n>0}\frac{p^{n}}{1-p^{n}}\alpha_{-n}z^{n}\right\}\exp\left\{\sum_{n>0}\frac{1}{1-p^{n}}\alpha_{n}z^{-n}\right\},$
$\hskip-28.45274pt\psi^{-}(\hbar^{-1/4}z)=\hbar^{-c/2}\exp\left\{-\sum_{n>0}\frac{1}{1-p^{n}}\alpha_{-n}z^{n}\right\}\exp\left\{\sum_{n>0}\frac{p^{n}}{1-p^{n}}\alpha_{n}z^{-n}\right\}.$
@@
2.4. The level-(0,-1) representation
We next consider the level-(0,-1) representation of $\mathcal{U}$} called the $q$-Fock representation. For $u\in{\mathbb{C}}^{\times}$, let ${\mathcal{F}}_{u}$ be a vector space spanned by ${|\lambda\rangle}_{u}$ $(\lambda\in{\mathcal{P}})$, where
$\hskip-28.45274pt{\mathcal{P}}=\left\{\lambda=(\lambda_{1},\lambda_{2},\cdots)\ \left|\ \begin{matrix}\lambda_{i}\geq\lambda_{i+1},\ \lambda_{i}\in{\mathbb{Z}}_{\geq 0},\cr\lambda_{\ell}=0\ \mbox{for sufficiently large}\ \ell\cr\end{matrix}\ \right.\right\}$.
We denote by $\ell(\lambda)$ the length of $\lambda\in{\mathcal{P}}$ i.e. $\lambda_{\ell(\lambda)}>0$ and $\lambda_{\ell(\lambda)+1}=0$. We also set $|\lambda|=\sum_{i\geq 1}\lambda_{i}$ and denote by $\lambda^{\prime}$ the conjugate of $\lambda$.
@@ 2.4. The level-(0,-1) representation
The following action gives a level-(0,-1) representation of $\mathcal{U}$} on ${\mathcal{F}}_{u}$. We denote this by ${\mathcal{F}}^{(0,-1)}_{u}$.
$\gamma^{1/2}{|\lambda\rangle}_{u}={|\lambda\rangle}_{u},$
$x^{+}(z){|\lambda\rangle}_{u}=a^{+}(p)\sum_{i=1}^{\ell(\lambda)+1}A^{+}_{\lambda,i}(p)\delta(u_{i}/z){|\lambda+{\bf 1}_{i}\rangle}_{u},$
$x^{-}(z){|\lambda\rangle}_{u}=a^{-}(p)\sum_{i=1}^{\ell(\lambda)}A^{-}_{\lambda,i}(p)\delta(t_{1}u_{i}/z){|\lambda-{\bf 1}_{i}\rangle}_{u},$
$\psi^{+}(z){|\lambda\rangle}_{u}=B^{+}_{\lambda}(u/z;p){|\lambda\rangle}_{u},$
$\psi^{-}(z){|\lambda\rangle}_{u}=B^{-}_{\lambda}(z/u;p){|\lambda\rangle}_{u},\quad$
@@
$a^{+}(p)$ $=$ $(1-t)\frac{(p\hbar;p)_{\infty}(p/t_{2};p)_{\infty}}{(p;p)_{\infty}(p/q;p)_{\infty}},$
$a^{-}(p)$ $=$ $(1-t^{-1})\frac{(p/\hbar;p)_{\infty}(pt_{2};p)_{\infty}}{(p;p)_{\infty}(pq;p)_{\infty}},$
$A^{+}_{\lambda,i}(p)$ $=$ $\prod_{j=1}^{i-1}\frac{\theta(t_{2}u_{i}/u_{j})\theta(\hbar^{-1}u_{i}/u_{j})}{\theta(t_{1}^{-1}u_{i}/u_{j})\theta(u_{i}/u_{j})},$
$A^{-}_{\lambda,i}(p)$ $=$ $\prod_{j=i+1}^{\ell(\lambda)}\frac{\theta(\hbar^{-1}u_{j}/u_{i})}{\theta(u_{j}/u_{i})}\prod_{j=i+1}^{\ell(\lambda)+1}\frac{\theta(t_{2}u_{j}/u_{i})}{\theta(t_{1}^{-1}u_{j}/u_{i})},$
$B^{+}_{\lambda}(u/z;p)$ $=$ $\prod_{i=1}^{\ell(\lambda)}\frac{\theta(t_{2}^{-1}u_{i}/z)}{\theta(t_{1}u_{i}/z)}\prod_{i=1}^{\ell(\lambda)+1}\frac{\theta(\hbar u_{i}/z)}{\theta(u_{i}/z)},$
$B^{-}_{\lambda}(z/u;p)$ $=$ $\prod_{i=1}^{\ell(\lambda)}\frac{\theta(t_{2}z/u_{i})}{\theta(t_{1}^{-1}z/u_{i})}\prod_{i=1}^{\ell(\lambda)+1}\frac{\theta(\hbar^{-1}z/u_{i})}{\theta(z/u_{i})}.$
@@
Remark. In [32] it is conjectured that the level-(0,-1) representation in Theorem 2.5 gives a geometric action of $\mathcal{U}$} on the equivariant elliptic cohomology of the Hilbert schemes $\bigoplus_{n}\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$ by identifying ${|\lambda\rangle}_{u}$ with the fixed point class $\lambda]$ in $\bigoplus_{n}\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$. The class $\lambda]$ with $|\lambda|=n$ should be realized in terms of the elliptic stable envelopes $\mathrm{Stab}_{\mathfrak{C}}$ for $\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$ constructed in [44] (see Sec.3.2 below) by
$[\lambda]=\sum_{\mu\in{\mathcal{P}}\atop|\mu|=n}\mathrm{Stab}^{-1}_{\mathfrak{C}}(\mu)|_{\lambda}\mathrm{Stab}_{\mathfrak{C}}(\mu)$
@@
3. Elliptic Stable Envelopes for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$
In this section we summarize basic facts on the instanton moduli space ${\mathcal{M}}(n,r)$ and the elliptic stable envelopes for the equivariant elliptic cohomology $\mathrm{E}_{T}({\mathcal{M}}(n,r))$ following [1, 44, 4].
3.1. The instanton moduli space ${\mathcal{M}}(n,r)$
Let ${\mathcal{M}}(n,r)$ be the moduli space of framed rank $r$ torsion free sheaves $\mathcal{S}$} on ${\mathbb{P}}^{2}$ with $c_{2}({\mathcal{S}})=n$. A framing of a sheaf $\mathcal{S}$} means a choice of isomorphism
${\mathcal{S}}|_{{\mathbb{P}}}\ \to\ {\mathcal{O}}^{\oplus r}_{L_{\infty}},$
where $L_{\infty}\subset{\mathbb{P}}^{2}$ is chosen as a line at infinity of a plane in ${\mathbb{P}}^{2}$. The space ${\mathcal{M}}(n,r)$ is isomorphic to the Nakajima quiver variety associated to the quiver consisting of one vertex and one loop with the vertex dimension $n$ and the framing dimension $r$, and known as the Atiyah-Drinfeld-Hitchin-Manin (ADHM) instanton moduli space. One has a natural action of $G=GL(r)\times GL(2)$ on ${\mathcal{M}}(n,r)$. Let $T$ be the maximal torus of $G$ and set